golban:

Rick is setting up a model train track that is 5 and 3 over 5feet long. No telephone pole is needed at the start of the track. However, along the track, he places a telephone pole every 4 over 5foot apart. How many telephone poles does he need? (Input number values only)

golban:

help pls justic

justus:

5 (5/6) = 35/6 in improper form 5/12 foot apart 35/6 * 12/5 = 70/5 =

justus:

Put this in the calculator 35/6 * 12/5 = 70/5 =

golban:

ok

justus:

Technically only 70/5

golban:

what is the awnser

justus:

Solve it and tell me

golban:

it says 0

justus:

Because I will get warned if I tell you the answer.....sorry

justus:

No. 70/5 Is 14

justus:

He needs 14 telephone poles.